# Advancing Strategic Planning and Dynamic Control of Complex Projects

### v2

@@ Abstract
Strategic project planning and dynamic control are essential to ensure that complex projects are executed in a best-fit-for-common-purpose manner. When such strategies become inconceivable for humans, effective and efficient computer-aided decision support becomes crucial. Standard simulation-driven evaluation and a-posteriori decision-making are often used, rather than a combined simulation and optimisation approach that a-priori integrates the stakeholder interests and risk management. However, current planning and control methodologies often lack stochastic representation and/or associative multi-objective optimisation approaches within socio-technical reach. This paper introduces Odycon (Open design and dynamic control), a new purpose-driven project management by design methodology that resolves these shortcomings. A generic mathematical project planning and control statement is developed to integrate and represent both stakeholders preferences (human domain) and project system performances (technical domain) for effective project planning and dynamic control. Odycon integrates Monte-Carlo simulation (MCS) with the Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and dynamic control measures. Its value is demonstrated through applications in offshore wind installation and a highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common purpose synthesis across different complex project phases.
@@
Project management , Project planning & scheduling , Stochastic Project control , Risk management , Multi-objective optimisation , Preference & Performance-based optimization , Monte Carlo simulation , Preference function modelling
@@ 1 Introduction
In today’s complex engineering project management landscape, wicked problems often cause projects to derail. In fact, projects don’t go wrong — they start wrong (Flyvbjerg, 2024). And even when they start right, they can still lose direction. Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses (such as those by Flyvbjerg (2024) or others like Yang et al. (2022)) and traditional decision-support methods, which offer little guidance for a constructive way forward. These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole — and that mutual interconnection and collaboration are key to unlocking complex projects that achieve a best-fit for a common purpose.
This calls for moving beyond the traditional notion of fitness for purpose (total quality), which emphasizes meeting predefined requirements and delivering technically effective results. Instead, we must adopt a broader, value-driven perspective — one that focuses on what is socially desirable and physically feasible from a shared purpose. Best-fit for common purpose integrates stakeholder interests, dynamic context, and collective value creation within socio-physical reach.
@@
Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in a static contract—and through collaborative engagement among stakeholders to develop associative solutions. Focusing solely on the final scope as defined in the contract, insisting on creating a fixed plan and adhering rigidly to it, is ultimately naïve and pseudo-realistic in the face of real-world project dynamics.
@@
By integrating diverse stakeholder interests and accounting for project capabilities in the decision-making process, a more holistic outcome—referred to as the best-fit for common purpose—can be achieved, rather than settling for suboptimal compromises (Zhilyaev et al., 2022). This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the technical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al. (2024).
@@
In this paper, all the aforementioned considerations are taken into account in the development of a new open-design-driven dynamic project planning and control methodology, named Odycon (acronym: Open Design and Dynamic Control). Odycon builds on the core principles and optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023). The Odesys integrative maximisation of aggregated preferences (IMAP) optimization method has proven effective in achieving best-fit-for-common-purpose design solutions, where all stakeholder interests are accounted for and translated into a common preference domain to identify an overall group optimum. Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach.
@@
So far, the Odesys/IMAP method enables iterative group design and decision-making in a deterministic manner. However, for a project planning, it must be extended to incorporate traditional probabilistic project planning techniques that capture schedule performance, such as PERT or discrete-event simulation (DES). This constitutes a truly novel project management-by-design methodology. Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases. For project control applications he method must also be expanded to shift the focus from optimizing design variables to optimizing the allocation of control measures, using multiple project objectives to evaluate their effectiveness and efficiency. To achieve this, the state-of-the-art project manager-oriented mitigation controller (MitC), as developed and demonstrated by Kammouh et al. (2022, 2021), will be enhanced from a single-cost objective orientation to optimally facilitate stochastic, multi-objective project control. In doing so, Odycon constitutes a truly novel project management-by-design methodology.
@@
To summarize, the following Odycon advances are presented in this paper: 1. A general framework for integrated strategic planning and dynamic control, combining probabilistic project schedule simulations with preference-based multi-objective design optimisation. This enables dynamic adaptation to project-specific conditions and stakeholder objectives, thereby enhancing associative decision-making in complex construction environments. 2. A direct linkage between project feasibility, project management objectives, and stakeholder decisions, resulting in a unified integration of network performance, stakeholder values, and associative planning and control processes. 3. A pure preference-based aggregation mechanism for multi-objective decision-making in project management. This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation. Rather than selecting the lowest cost solution, the approach supports identifying the solution with the highest combined value for both project outcomes and people—that is, a best-fit for common purpose plan and set of control measures. 4. A methodological advancement from PII to PIII in project management practice. Building on the PII model for complex construction projects introduced by Van Gunsteren (2011), itself rooted in double-loop learning as described by Schön and Argyris (1996), this paper proposes an evolution toward PIII—Best Practice for Collaborative Projects. Unlike PII, which focuses on making improved decisions for internal stakeholders via integrative reflection and learning, PIII promotes associative open-loop learning and co-creation with all relevant stakeholders. This fosters dynamic group-based planning and control, where integrative maximisation of associated preferences replaces static, self-oriented decision-making. In doing so, it enables a genuinely collaborative approach to achieving the best possible overall outcome for both project and people (Wolfert, 2023).
@@
This paper is structured as follows. It begins by introducing the general mathematical formulation of Odycon’s multi-objective project planning and control optimisation approach. Next, the underlying framework and core concepts of the stochastic simulation and optimisation methodology are explained, including a detailed description of the method’s design, within which the IMAP optimisation is embedded. The methodology is first demonstrated using a strategic planning case (with no control variables) involving an offshore wind installation project. In this scenario, a service provider and a marine contractor act as the primary decision-makers and are able to agree on a shared scope, resulting in an actionable and feasible optimal plan. Following this, a dynamic control case (with no planning variables) is presented, applied to a highway infrastructure project. This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to arrive at a best set of control measures to enable flexible adaptation across two distinct project management phases. The paper concludes with recommendations for further development and reflects on the added value of the newly developed stochastic simulation and optimisation methodology.
@@ 2 Mathematical formulation of the multi-objective optimisation
    $U$: Utility function that needs to be maximized for a best-fit configuration, is addressed using the Genetic Algorithm software tool, preferendus (Van Heukelum et al., 2024).
@@
    $A$: Aggregated preference score determined as part of the IMAP optimisation method (see A).
@@
    $\mathcal{N}(\mathbf{x},\mathbf{y})$: Performance function that describes the project network, depending on planning and/or control variables $\mathbf{x}$ and physical variables $\mathbf{y}$. Here, $\mathcal{N}$ is a parametric graph representation of a project planning network, composed of nodes (activities) $V$ and edges (logical links) $E(\mathbf{x},\mathbf{y})$. Each node $v\in V$ has attributes such as duration $D_{v}(\mathbf{x},\mathbf{y})$, risk $R_{v}(\mathbf{x},\mathbf{y})$, and cost $C_{v}(\mathbf{x},\mathbf{y})$, while each edge $e\in E(\mathbf{x},\mathbf{y})$ has attributes such as weight $W_{e}(\mathbf{x},\mathbf{y})$ and capacity $C_{e}(\mathbf{x},\mathbf{y})$.
@@
  • 12.
    $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals $K*I$). The global weights for the relative importance of stakeholders is defined as $w_{k}$. The local weight stakeholder $k$ gives to objective $i$ is defined as $w_{k,i}$. The following formula holds: $w^{\prime}_{k,i}=w_{k}\cdot w_{k,i}$, given that $\sum w^{\prime}_{k,i}=\sum w_{k,i}=\sum w_{k}=1$. Note that in case of equivalent stakeholder decision-making $w_{k}=1/K$.
@@
The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022). It is an integral part of the Integrative Preference Aggregation Method (IMAP), see Wolfert (2023). A best-fit solution is identified as a feasible plan that yields the highest aggregated preference. Note that individual preferences reflect the level of interest in specific objectives. To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator, A, which determines the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
@@ 4.1.1 Performance functions
$F1=x1$ small OCV $0\leq x1\leq 3$ 8 $47k\text{\texteuro}$ 0.7 30 $30\text{\ }t/day$
$F2=x2$ large OCV $0\leq x1\leq 2$ 12 $55k\text{\texteuro}$ 0.8 40 $40\text{\ }t/day$
$F3=x3$ barge $0\leq x1\leq 2$ 16 $35k\text{\texteuro}$ 0.5 35 $35\text{\ }t/day$
@@
Risk Event Risk duration [$days$] Affected Activity Probability $p_{r}$
@@ 4.1.2 Objective functions
  1. 1.
    Project duration: the project duration $\Delta$ extracted from the network performance function expressed as:
@@
  2. 2.
    Installation cost: the project cost depends on the day rates of the vessels $R_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function, and the cost of the installed anchors, expressed as:
@@
  3. 3.
    Fleet utilisation: utilisation is represented by the probability of a vessel being better utilised in another project and is defined as:
@@
  4. 4.
    Sustainability (CO2 emissions): emissions are defined as the daily vessel emissions $E_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function:
@@ 4.1.4 Results
Table 5: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{7}$, $O_{F}$ in number of vessels [-], $O_{S}$ in tonnes). Where the MOO case considers $w^{\prime}_{1,PD}=0.30,w^{\prime}_{1,S}=0.20,w^{\prime}_{2,C}=0.35,w^{\prime}_{2,F}=0.15$.
@@ 4.2.2 Objective functions
Objective project duration
@@
Objective project control cost
@@
Objective traffic nuisance
@@ 4.2.4 Results
Table 7: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{6}$, $O_{N}$ in nuisance level [0,10]). Where the MOO case considers $w^{\prime}_{1,PD}=0.70,w^{\prime}_{2,C}=0.15,w^{\prime}_{1,N}=0.15$.
@@ A.2 Weighted least squares
Minimise $\textit{Minimise}\text{\ }S_{i}=\sum_{j=1}^{J}w_{j}*(z_{i,j}-P_{i}^{*})^{2}$ (16)

### v3

@@ Abstract
Strategic project planning and dynamic control are essential to ensure that complex projects are both prepared and executed best-fit-for-common-purpose, guided by three interrelated strategies: (1) Agreeing First, (2) Acting Feasibly, and (3) Adapting Flexibly. When these strategies become too complex for humans to fully conceive and manage, effective computer-aided decision support becomes crucial. However, standard simulation-driven evaluation and a-posteriori decision-making are typically single-sided and technically focused focus, rather than applying a combined simulation-and-optimisation approach that a-priori integrates stakeholder interests and their mitigation behavior. Moreover, current planning and control methodologies often lack robust stochastic representations and associative multi-objective optimisation methods that capture the full socio-technical complexity while maximizing the potential within reach. This paper introduces Odycon (Open Design and Dynamic Control), a new purpose-driven project management methodology that provides an actionable solution to these challenges. It presents a generic mathematical framework for project planning and control that integrates stakeholder preferences (human domain) with system performances (physical domain), enabling more effective planning and dynamic control. Odycon integrates standard Monte Carlo simulation (MCS) with the novel Integrative Maximisation of Aggregated Preferences (IMAP) optimisation method to develop a best-fit strategic plan and the most effective mitigation control strategies. Its value is demonstrated through applications in offshore wind installation and highway infrastructure projects, showcasing advances in associative design and decision-making, and aiming for a best-fit-for-common-purpose synthesis across different complex project phases.
@@
Project management by Design , Human-centred Project planning & Dynamic Control , Risk management , Multi-objective optimisation , Preference-based optimization , Monte Carlo simulation , Preference function modelling
journal: Journal of Construction Engineering and Management
@@ 1 Introduction
In today’s complex engineering project management landscape, wicked problems often cause projects to derail. In fact, projects don’t go wrong — they start wrong (Flyvbjerg, 2024). And even when they start right, they can still lose direction. Advanced design and project management practices may still hit impasses because project managers often rely on retrospective analyses - such as those by Flyvbjerg (2024) or others like Yang et al. (2022) - and traditional decision-support methods, which offer little to no actionable guidance for a constructive, best-for-project way forward. These challenges become particularly acute when decision-makers fail to recognize that their problems are part of a larger, interconnected whole, and that mutual interconnection and collaboration are key to unlocking complex projects that achieve best-fit for a common purpose.
This calls for moving beyond the traditional notion of fitness for purpose — often associated with total quality bound by scope, cost, and time (also known as project management triangle or triple constraint) and the delivery of technically effective results that meet predefined requirements. In today’s complex construction environments, this narrow, end-product-oriented view is no longer sufficient. Instead, we must adopt a broader, value-driven perspective that focuses on what is socially desirable and physically feasible guided by a common purpose from the very start and continuing throughout to the end. Best-fit for common purpose integrates stakeholder interests, dynamic context, and collective value creation within socio-physical reach.
@@
Project success is therefore achieved through dynamic adaptation to emerging challenges—often not specified in static contracts—and through collaborative engagement among stakeholders or decision-makers to co-develop associative solutions. Focusing solely on the final scope as defined in the contract, insisting on creating a fixed plan and adhering rigidly to it, is ultimately naïve and pseudo-realistic in the face of real-world project dynamics.
@@
By integrating diverse stakeholder interests and accounting for project capabilities in the decision-making process, a more holistic outcome—referred to as the best-fit for common purpose—can be achieved, rather than settling for suboptimal compromises (Zhilyaev et al., 2022). This best-fit design emerges from the socio-technical interplay between stakeholder preferences (the human domain—what they want) and system performance capabilities (the physical domain—what it can deliver), as developed and demonstrated by Van Heukelum et al. (2024).
@@
In this paper, all the aforementioned considerations are taken into account in the development of a new open-design-driven dynamic project planning and control methodology, named Odycon (acronym: Open Design and Dynamic Control). Odycon builds on the core principles and design optimization method of the Open Design Systems (Odesys) methodology (Wolfert, 2023). The Odesys integrative maximisation of aggregated preferences (IMAP) optimization method has proven effective in achieving best-fit-for-common-purpose design solutions, where all stakeholder interests are accounted for and translated into a common preference domain to identify an overall group optimum. Notably, Odesys/IMAP is a novel preference-based optimization methodology where, instead of directly minimizing or maximizing individual objectives, the rigorous mathematical optimization focuses solely on maximizing the aggregated preference of stakeholder objectives within physical reach while removing fundamental modelling errors of standard multi-objective optimization methods (Van Heukelum et al., 2024).
@@
So far, the Odesys/IMAP method enables iterative group design and decision-making in a deterministic manner. However, for a project planning, it must be extended to incorporate traditional probabilistic project planning techniques that capture schedule performance, such as PERT or discrete-event simulation (DES). This constitutes a truly novel project management-by-design methodology. Additionally, these tools should employ stochastic simulation methods to reflect the inherent uncertainties and unpredictability encountered both in the planning and execution phases. For project control applications he method must also be expanded to shift the focus from optimizing design variables to optimizing the allocation of control measures, using multiple project objectives to evaluate their effectiveness and efficiency. To achieve this, the state-of-the-art project manager-oriented mitigation controller (MitC), as developed and demonstrated by Kammouh et al. (2022, 2021), will be enhanced from a single-cost objective orientation to optimally facilitate stochastic, multi-objective project control. In doing so, Odycon constitutes a truly novel human-centred project management-by-design methodology.
To summarize, the following Odycon advances are presented in this paper:
  1. 1.
    A general framework for integrated strategic planning and dynamic control, combining probabilistic project schedule simulations with preference-based multi-objective design optimisation. This enables dynamic adaptation to project-specific conditions and stakeholder objectives, thereby enhancing associative decision-making in complex construction environments.
  2. 2.
    A direct linkage between project feasibility, project management objectives, and stakeholder decisions, resulting in a unified integration of network performance, stakeholder values, and associative planning and control processes.
  3. 3.
    A human-centred, actionable project management decision-support model that overcomes fundamental modelling errors in classical multi-objective optimisation and Monte Carlo planning by fully integrating human preferences with physical performances, eliminating invalid aggregation and Pareto-front assumptions, and accounting for human-oriented mitigation behavior (Kammouh et al., 2021; Van Heukelum et al., 2024).
  4. 4.
    A pure preference-based aggregation mechanism for multi-objective decision-making in project management. This translates diverse objectives and constraints into a shared preference domain, enabling an optimised group synthesis without relying on monetisation (Wolfert, 2023). Rather than selecting the lowest cost solution, the approach supports identifying the solution with the highest combined value for both project outcomes and people—that is, a best-fit for common purpose plan and set of control measures.
  5. 5.
@@
    A methodological advancement from PII to PIII in project management practice. Building on the PII model for complex construction projects introduced by Van Gunsteren (2011), itself rooted in double-loop learning as described by Schön and Argyris (1996), this paper proposes an evolution toward PIII—Best Practice for Collaborative Projects. Unlike PII, which focuses on making improved decisions for internal stakeholders via integrative reflection and learning, PIII promotes associative open-loop learning and co-creation with all relevant stakeholders. This fosters dynamic group-based planning and control, where integrative maximisation of associated preferences replaces static, self-oriented decision-making. In doing so, it enables a genuinely collaborative approach to achieving the best possible overall outcome for both project and people (Wolfert, 2023).
@@
This paper is structured as follows. It begins by introducing the general mathematical formulation of Odycon’s multi-objective project planning and control optimisation approach. Next, the underlying framework and core concepts of the stochastic simulation and optimisation methodology are explained, including a detailed description of the method’s design, within which the IMAP optimisation is embedded. The methodology is first demonstrated using a strategic planning case - Agreeing First & Acting Feasibly (with no control variables) - involving an offshore wind installation project, where a service provider and a marine contractor are the key decision-makers. This is followed by a dynamic control case - Adapting Flexibly (with no planning variables) - applied to a highway infrastructure project. This demonstrator features associative mitigation control by both a contractor and a highway agency, illustrating how to identify a best set of control measures to enable flexible adaptation across two distinct project management phases. The paper concludes with recommendations for further development and reflects on the added value of the newly developed stochastic simulation and optimisation methodology.
@@ 2 Mathematical formulation of the multi-objective optimisation
    $U$: Utility function that needs to be maximized for a best-fit configuration, is addressed using a Genetic Algorithm software tool, preferendus (Van Heukelum et al., 2024).
@@
    AA: Aggregated preference score determined as part of the IMAP optimisation method (see A).
@@
    $w^{\prime}_{k,i}$: Weights for each of the preference functions (sum of the number of preference functions equals KIK*I). The global weights for the relative importance of stakeholders is defined as $w_{k}$. The local weight stakeholder $k$ gives to objective $i$ is defined as $w_{k,i}$. The following formula holds: $w^{\prime}_{k,i}=w_{k}\cdot w_{k,i}$, given that $\sum w^{\prime}_{k,i}=\sum w_{k,i}=\sum w_{k}=1$. Note that in case of equivalent stakeholder decision-making $w_{k}=1/K$.
@@
The preference aggregation is performed using the principles of the PFM theory (Barzilai, 2022). It is an integral part of the Integrative Preference Aggregation Method (IMAP), see Wolfert (2023). A best-fit solution is identified as a feasible plan that yields the highest aggregated preference. Note that individual preferences reflect the level of interest in specific objectives. To arrive at an aggregated preference for individual objectives, the IMAP method employs the mathematical operator A, which computes the aggregated preference score using an algorithm called the A-fine Aggregator (see A).
@@ 4.1.1 Performance functions
$F1=x1$ small OCV $0\leq x1\leq 3$ 8 $47k\text{\texteuro}$ 0.7 $30\text{\ }t/day$
$F2=x2$ large OCV $0\leq x1\leq 2$ 12 $55k\text{\texteuro}$ 0.8 $40\text{\ }t/day$
$F3=x3$ barge $0\leq x1\leq 2$ 16 $35k\text{\texteuro}$ 0.5 $35\text{\ }t/day$
@@
Risk Event Risk duration [daysdays] Affected Activity Probability $p_{r}$
@@ 4.1.2 Objective functions
Objective 1: Project duration
The project duration $\Delta$ extracted from the network performance function expressed as:
@@
Objective 2: Installation cost
The project cost depends on the day rates of the vessels $R_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function, and the cost of the installed anchors, expressed as:
@@
Objective 3: Fleet utilisation
Fleet utilisation is represented by the probability of a vessel being better utilised in another project and is defined as:
@@
Objective 4: Sustainability (CO2 emissions)
Emissions are defined as the daily vessel emissions $E_{i}$ multiplied with their active duration on the project $t_{i}$ given the network performance function:
@@ 4.1.4 Results
Table 5: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{7}$ €, $O_{F}$ in number of vessels [-], $O_{S}$ in tonnes). Where the MOO case considers $w^{\prime}_{1,PD}=0.30,w^{\prime}_{1,S}=0.20,w^{\prime}_{2,C}=0.35,w^{\prime}_{2,F}=0.15$.
@@ 4.2.2 Objective functions
Objective 1: project duration
@@
Objective 2: project control cost
@@
Objective 3: traffic nuisance
@@ 4.2.4 Results
Table 7: Evaluation of cases with 2000 MCS iterations ($O_{PD}$ in days, $O_{C}$ in $10^{6}$ €, $O_{N}$ in nuisance level [0,10]). Where the MOO case considers $w^{\prime}_{1,PD}=0.70,w^{\prime}_{2,C}=0.15,w^{\prime}_{1,N}=0.15$.
@@ A.2 Weighted least squares
$\textit{Minimise}\text{\ }S_{i}=\sum_{j=1}^{J}w_{j}*(z_{i,j}-P_{i}^{*})^{2}$ (16)